The risk-free rate is 2.74% and the market risk premium is 5.27%. A stock with a β of 1.02 just paid a dividend of $1.92. The dividend is expected to grow at 21.81% for three years and then grow at 4.37% forever. What is the value of the stock?
Answer #1

Cost of equity = Risk free rate + Beta * market risk premium= 2.74% + 1.02* 5.27% = 8.1154%= approx 8.12%

Current dividend, D0 = $1.92

Dividend next year, D1 = 1.92* (1+ 21.81%) = 2.34

D2 = 2.24*(1.2181) = $ 2.85

D3 = 2.85*1.2181 = .$ 3.47

Terminal Value of dividend = D3*(1+g)/(k-g) = 3.47*(1+4.37%)/(8.12% - 4.37%)

= $ 96.58

Value of stock = 2.34/(1+8.12%) + 2.85/(1+8.12%)^2 + 3.47/(1+8.12%)^3 + 96.58/(1+8.12%)^3

= $ 85.34